If I have 30 liters of a 10% solution, how many liters of a 2% solution do I add to get an 8% soolution?
Answer by amarjeeth123(569)

Let the number of litres be x.
We have 30(10)+x(2)=(x+30)8
300+2x=8x+240
6x=60
x=10
10 litres of a 2% solution must be added.